Solution A is 10% sugar while Solution B is 30% sugar.A chemist wants to have a 60-liter mixture of solutions A and B with 18% sugar.How many liters of each solution should he use?
Answer by stanbon(75887)

Solution A is 10% sugar while Solution B is 30% sugar.A chemist wants to have a 60-liter mixture of solutions A and B with 18% sugar.How many liters of each solution should he use?
------------------------
Equation:
sugar + sugar = sugar
0.10x + 0.30(60-x) = 0.18*60
---
Multiply thru by 100 to get:
10x + 30*60 - 30x = 18*60
-20x = -12*60
x = 36 liters (amt. of 10% solution needed in the mixture)
---
60-x = 24 liters (amt of 0.30% solution needed in the mixture)
====================
Cheers,
Stan H.